OpenStudy (anonymous):

Any online calculators to shorten stat->edit? Are there any calculator that can basically shorten the Stat->Edit mess of having to input every x-value? Lets say I want to input every number from 280 to 1600 that is divisible by 5. Is there anything other than a algorithm I couldn't code myself that I can just set an upperbound (265), and use the formula 5(n+55) , from n=1 to 265 to input values?

OpenStudy (amistre64):

L1 = seq(fcn, var, first, last, incrent) L1 = seq(X,X,280,1600,1) should populate L1 with 280 thru 1600

OpenStudy (amistre64):

spose you had some function (2x^2) that you wanted to populate L2 with from L1 L2 = seq(2*(L(X))^2, X, 0,range, 1)